# On Lagrangian formulations for (ir)reducible mixed-antisymmetric higher integer spin fields in Minkowski spaces

### v1

@@ 1 Introduction
Problems of higher-spin (HS) field theory have been the subject of significant discussions among theoretician community. Interest in HS field theory (for review, see e.g. [1], [2], [3], [4]) is mainly inspired by hope to develop the new Lagrangian models in classical field theory and may be formulate on this ground the new tools for the unification of the fundamental interactions. HS field theory is closely related to (super)string theory [5], [6] which operates with an infinite tower of bosonic and fermionic HS fields in $10$-dimensional space-time with compactified extra dimensions thus generating theirs masses within dimensional reduction mechanism. Because of interacting (super)strings seems to be only unique way to consistently describe evolution of early Universe and involve the interaction of particles with different values of spin, starting from known particles with lower spins (Higgs particle for $s=0$; leptons and quarks for $s=\frac{1}{2}$; photon $W^{\pm},Z$ bosons for $\lambda=1$ and yet undiscovered graviton for $\lambda=2$) the unknown particles both with HS $s>2$ and with generalized spin $s_{(k)}=(s_{1},s_{2},...,s_{k})$ are also admissible. From this perspective such particles may be considered as alternative candidate to describe Dark Matter, see for a review[7], [8] (always surrounding electromagnetically observable particles) beyond the extensions of the Standard Model involving sterile neutrino [9] and dark massive photon [10]. Whereas the problem of LF construction for HS fields in constant curvature spaces have been considered in different approaches (see for review, [11], [12]) both in metric and frame-like formalism for totally, e.g. in [13], [14], [15], [16], [17], [18] and mixed-symmetric irreducible (spin)-tensors fields on $\mathbb{R}^{1,d-1}$ (or AdS space), e.g. in [19], [20], [21], [22], [23], [24], [25], [26], [27] (for $d=4$ as well as for irreducible representations (irreps) with continuous spin, see [28], [29], [30], [31]), the consistent interacting LF with HS fields is still far from its complete realization, in spite for the results (see, e.g., the papers and references therein) for cubic [33], [32], [34], [35], [36], [37], [38] and quartic [39], [40] vertices.
@@
The first purpose of the paper is to construct general gauge-invariant Lagrangians for both massless and massive mixed-symmetry tensor fields of rank $s_{1}+s_{2}+s_{3}$, with any integer numbers $s_{1}\geq s_{2}\geq s_{3}\geq 1$ for $s_{1}<[d/2]$ in $\mathbb{R}^{1.d-1}$, as elements of Poincare-group ISO(1,d1)ISO(1,d-1) irreps with a Young tableau having $3$ columns. Another aim is to formulate the deformation procedure to construct interacting gauge model with mixed-antisymmetric fields.
@@ 2 Lagrangian formulation with complete BRST operator \\
Here, we shortly recall [48] that a massless (massive) integer spin irreducible representation of Poincare group in a Minkowski space $\mathbb{R}^{1,d-1}$ which is described by a real-valued tensor field $\Phi_{\mu^{1}[{\hat{s}_{1}}],\mu^{2}[{\hat{s}_{2}}],\mu^{3}[{\hat{s}_{3}}]}\hskip-1.99997pt$ $\equiv$ $\hskip-1.99997pt\Phi_{\mu^{1}_{1}\ldots\mu^{1}_{\hat{s}_{1}},...,\mu^{3}_{1}\ldots\mu^{3}_{\hat{s}_{3}}}$ of rank $\sum_{i}^{3}\hat{s}_{i}$ and generalized spin $\mathbf{s}=(3,3,...,3,2,2,...,2,1$ ,,1),...,1), ($\hat{s}_{1}\geq\hat{s}_{2}\geq s_{3}>0,\hat{s}_{1}\leq[d/2])$ to be corresponding to a Young tableau with $3$ columns of height $\hat{s}_{1},\hat{s}_{2},\hat{s}_{3}$, respectively (with omitting the symbol $\hat{\phantom{s}}"$ under $\hat{s}_{i}$)
@@
It is easy to see, the system (3)–(6) is equivalent to set of primary constraints on the string=like vector $|\Phi\rangle\in\mathcal{H}^{f}$ in an auxiliary Fock space $\mathcal{H}^{f}$, generated by $3$ pairs of Grassmann-odd (antisymmetric basis) oscillators $\hat{a}^{i+}_{\mu^{i}}(x)$, $\hat{a}^{j}_{\nu^{j}}(x)$, $\mu^{i},\nu^{j}=0,1...,d-1$:
@@
$\hskip-8.61108pt\bigl{(}{l}_{0},\,{l}^{i},\,l^{ij},\,t^{ij},\,g_{0}^{i}-(s_{i}-\textstyle\frac{d}{2})\bigr{)}|\Phi\rangle=\bigl{(}\partial^{\mu}\partial_{\mu},\,-i\hat{a}^{i}_{\mu}\partial^{\mu},\,\textstyle\frac{1}{2}\hat{a}^{i}_{\mu}\hat{a}^{j\mu},\,\hat{a}^{i+}_{\mu}\hat{a}^{j\mu},\,-\hat{a}^{i+}_{\mu}\hat{a}^{\mu{}i}-\frac{d}{2}\bigr{)}|\Phi\rangle=\vec{0}.$ (8)
@@
The operators $o^{\prime}_{I}$ are hermitian with respect to modified (by operator $K^{\prime}$ intertwining $\{|\vec{N}\rangle_{V}\}$ and $\{|\vec{N}\rangle\}$ (32)) scalar product. Having in mind the rules of the dimensional reduction from massless theory in $\mathbb{R}^{1,d}$ space-time to massive one in $\mathbb{R}^{1,d-1}$ with fermionic oscillators $f_{i}^{+},f_{i}$,
$\partial^{M}=(\partial^{\mu},-\imath m)\,$, $\hat{a}^{M}_{i}=(\hat{a}^{\mu}_{i},f_{i})\,$, $\hat{a}^{M{}+}_{i}=(\hat{a}^{\mu{}+}_{i},f_{i}^{+})\,$, (21)
$M=0,1,\ldots,d\,$, $\mu=0,1,\ldots,d-1\,$, $\eta^{MN}=diag(1,-1,\ldots,-1,-1)\,$, (22)
the converted operators $O_{I}$ now for massive case are derived as well.
@@ 2.1 BRST operator and Lagrangian formulation
(for $gh_{H}(\mathcal{C}^{I})=-gh_{H}(\mathcal{P}_{I})=1$). Here, the Grassman-even ghost coordinates $q_{i}^{+}$, $q_{i}$ momenta $p_{j}$, $p_{j}^{+}$ for Grassman-odd basis elements $l_{i}$, $l_{i}^{+}$ of the superalgebra $\mathcal{A}_{C}(Y[k],\mathbb{R}^{1,d-1})$, respectively, and Grassman-odd ghost coordinates $\eta_{0}$, $\eta_{ij}^{+}$, $\eta_{ij}$, $\vartheta_{rs}^{+}$, $\vartheta_{rs}$, $\eta_{i}^{G}$ and momenta $\mathcal{P}_{0}$, $\mathcal{P}_{ij}$, $\mathcal{P}_{ij}^{+}$, $\lambda_{rs}$, $\lambda_{rs}^{+}$, $P^{G}_{i}$ for Grassman-even basis elements $l_{0}$, $L_{ij}$, $L_{ij}^{+}$, $T_{rs}$, $T_{rs}^{+}$, $G_{0}^{i}$ of $\mathcal{A}_{C}(Y[k],\mathbb{R}^{1,d-1})$, respectively, are introduced subject to non-trivial commutation relations,
@@
it follows a finite sequence of relations underlying the $\eta^{i}_{G}$-independence of all of the above homogeneous in ghost number vectors:
$Q|\chi\rangle=0,$ $(\sigma^{i}+h_{i})|\chi\rangle=0,$ $\left(\varepsilon,{gh}_{H}\right)(|\chi\rangle)=(\varepsilon_{\chi},0),$ (35)
$\delta|\chi\rangle=Q|\chi^{1}\rangle,$ $(\sigma^{i}+h_{i})|\chi^{1}\rangle=0,$ $\left(\varepsilon,{gh}_{H}\right)(|\chi^{1}\rangle)=(\varepsilon_{\chi}+1,-1),$ (36)
$\ldots\ldots$ $\ldots\ldots$ $\ldots\ldots$
$\delta|\chi^{n-1}\rangle=Q|\chi^{n}\rangle,$ $(\sigma^{i}+h_{i})|\chi^{n}\rangle=0,$ $\left(\varepsilon,{gh}_{H}\right)(|\chi^{n}\rangle)=(\varepsilon_{\chi}+n\hskip-5.0pt\mod{}2,-n).$ (37)
By one of the peculiarity of BRST approach to mixed-antisymmetric HS fields is the fact that the Grassmann parity of any of $|\chi^{n}\rangle$ depends on the values of spin. The solutions of middle set of equations in (35)–(37) determine values of spin for the eigenvectors $\{|\chi^{n}\rangle_{{s[3]}}\}$ of generalized spin operator $(\sigma^{1},\sigma^{2},\sigma^{3})$ and possible values for parameters $h_{i}$:
@@
As the result the operator $Q$ is nilpotent on the solutions for the middle set and the first equation in (35): $Q|\chi\rangle_{{s[3]}}=0$ is Lagrangian EoM and follows from gauge-invariant LF
@@
(for substitution $(K_{s[3]},Q_{s[3]})=(K,Q)|_{h^{i}=h^{i}(s[3])}$) with reducible gauge transformations
$\delta|\chi^{0}\rangle_{s[3]}=Q_{s[3]}|\chi^{1}\rangle_{s[k]}\,,\ldots,\delta|\chi^{l}\rangle_{s[3]}=Q_{s[k]}|\chi^{l+1}\rangle_{s[3]},\ l=\sum_{i=1}^{3}s_{i}+2\,$, (40)
for independent gauge parameter $|\chi^{l+1}\rangle$. Here the usual inner product for the creation and annihilation operators is assumed with measure $d^{d}x$ over Minkowski space with additional terms: "more" with auxiliary fields differed for massless and massive basic HS field $\Phi^{\mu[{s_{1}}],\nu[{s_{2}}],\rho[{s_{3}}]}(x)$.
@@ 3 Lagrangian formulation with incomplete BRST operator \\
The set of $Q_{c},\sigma^{i}_{c},\widehat{L}_{ij},\widehat{T}_{ij}$ acts in Hilbert space $\mathcal{H}^{f}\otimes H_{gh}^{o_{A}}$ and forms closed superalgebra/ It provides the consistent BRST $Q_{c}$-complex resolution to get LF with holonomic operator constraints imposed on whole set of restricted field $|\chi_{c}\rangle_{s[3]}$ and reducible gauge parameter $|\chi_{c}^{n}\rangle_{s[3]}$ vectors. The latter vectors have the decomposition in oscillators like (2.1) but without conversion $b^{(+)}_{ij},d^{(+)}_{rs}$ and ghost oscillators $\eta_{ij}^{+},\mathcal{P}_{ij}^{+},\vartheta_{rs}^{+},\lambda_{rs}^{+}$ for second-class constraint system.
@@ 4 Deformation procedure for interacting vertices \\
To this end we introduce $p$, $p\geq 3$, copies of LFs (in adapting the model for Yang-Mills type interactions with gauge group $SU(N)$, for $p=N^{2}-1$) with vectors $|\chi^{(t)}_{c}\rangle_{{s}[k]_{t}}$, reducible gauge parameters $|\Lambda^{(t)n}_{c}\rangle_{{s}[k]_{t}}\equiv|\chi^{(t)n+1}_{c}\rangle_{{s}[k]_{t}}$, respective vacuum vectors $|0\rangle^{t}$ and oscillators for t=1,,pt=1,...,p (with notation $(s[3])_{t}\equiv(s[3]_{1},...,s[3]_{t})$ for arbitrary sets of indices for different copies). It permits one to determine the deformed action, gauge transformations up to $r$-tic vertices, r=3,4,,er=3,4,...,e in powers of deformation constant $g$, beginning from sum of $p$ LFs copies for free HS fields, then for cubic, quartic and so on vertices:
@@
$\hskip-5.0pt\delta^{l}_{e}|\Lambda^{(t)l}_{c}\rangle_{{s}[3]_{t}}=-\sum_{1\leq i_{1}<...<i_{e-1}\leq p}\hskip-10.00002pt\int\prod_{j=1}^{e-1}d\eta^{(i_{j})}_{0}\Big{[}{}_{{s}[3]_{{i_{1}}}}\langle\chi^{(\{i_{1})}_{c}\big{|}\ldots{}_{{s}[3]_{{i_{e-2}}}}\langle\chi^{(i_{e-2})}_{c}\big{|}$ (49)
@@
The resolution of (50) for xx=local cubic vertices with representation

### v2

@@ 1 Introduction
Problems of higher-spin (HS) field theory have been the subject of significant discussions among theoretician community. Interest in HS field theory (for review, see e.g. [1], [2], [3], [4]) is mainly inspired by hope to develop the new Lagrangian models in classical field theory and may be formulate on this ground the new tools for the unification of the fundamental interactions. HS field theory is closely related to (super)string theory [5], [6] which operates with an infinite tower of bosonic and fermionic HS fields in $10$-dimensional space-time with compactified extra dimensions thus generating theirs masses within dimensional reduction mechanism. Because of interacting (super)strings seems to be only unique way to consistently describe evolution of early Universe and involve the interaction of particles with different values of spin, starting from known particles with lower spins (Higgs particle for $s=0$; leptons and quarks for $s=\frac{1}{2}$; photon, $W^{\pm},Z$ bosons for $\lambda=1$ and yet undiscovered graviton for $\lambda=2$) the unknown particles both with HS $s>2$ and with generalized spin $s_{(k)}=(s_{1},s_{2},...,s_{k})$ are also admissible. From this perspective such particles may be considered as alternative candidate to describe Dark Matter, see for a review[7], [8] (always surrounding electromagnetically observable particles) beyond the extensions of the Standard Model involving sterile neutrino [9] and dark massive photon [10]. Whereas the problem of Lagrangian formulation (LF) construction for HS fields in constant curvature spaces have been considered in different approaches (see for review, [11], [12]) both in metric and frame-like formalism for totally, e.g. in [13], [14], [15], [16], [17], [18] and mixed-symmetric irreducible (spin)-tensors fields on $\mathbb{R}^{1,d-1}$ (or AdS space), e.g. in [19], [20], [21], [22], [23], [24], [25], [26], [27] (for $d=4$ as well as for irreducible representations (irreps) with continuous spin, see [28], [29], [30], [31]), the consistent interacting LF with HS fields is still far from its complete realization, in spite for the results (see, e.g., the papers and references therein) for cubic [33], [32], [34], [35], [36], [37], [38] and quartic [39], [40] vertices.
@@
The first purpose of the paper is to construct general gauge-invariant Lagrangians for both massless and massive mixed-symmetry tensor fields of rank $\hat{s}_{1}+\hat{s}_{2}+\hat{s}_{3}$, with any integer numbers $\hat{s}_{1}\geq\hat{s}_{2}\geq\hat{s}_{3}\geq 1$ for $\hat{s}_{1}<[d/2]$ in $\mathbb{R}^{1.d-1}$, as elements of Poincare-group ISO(1,d1)ISO(1,d-1) irreps with a Young tableau having $3$ columns. Another aim is to formulate the deformation procedure to construct interacting gauge model with mixed-antisymmetric fields.
@@ 2 Lagrangian formulation with complete BRST operator \\
Here, we shortly recall [48] that a massless (massive) integer spin irreducible representation of Poincare group in a Minkowski space $\mathbb{R}^{1,d-1}$ which is described by a real-valued tensor field $\Phi_{\mu^{1}[{\hat{s}_{1}}],\mu^{2}[{\hat{s}_{2}}],\mu^{3}[{\hat{s}_{3}}]}\hskip-1.99997pt$ $\equiv$ $\hskip-1.99997pt\Phi_{\mu^{1}_{1}\ldots\mu^{1}_{\hat{s}_{1}},...,\mu^{3}_{1}\ldots\mu^{3}_{\hat{s}_{3}}}$ of rank $\sum_{i}^{3}\hat{s}_{i}$ and generalized spin $\mathbf{s}=(3,3,...,3,2,2,...,2,1$ ,,1),...,1), ($\hat{s}_{1}\geq\hat{s}_{2}\geq\hat{s}_{3}>0,\hat{s}_{1}\leq[d/2])$ to be corresponding to a Young tableau with $3$ columns of height $\hat{s}_{1},\hat{s}_{2},\hat{s}_{3}$, respectively (with omitting further the symbol $\hat{\phantom{s}}"$ under $\hat{s}_{i}$)
@@
It is easy to see, the system (3)–(6) is equivalent to set of primary constraints on the string–like vector $|\Phi\rangle\in\mathcal{H}^{f}$ in an auxiliary Fock space $\mathcal{H}^{f}$, generated by $3$ pairs of Grassmann-odd (antisymmetric basis) oscillators $\hat{a}^{i+}_{\mu^{i}}(x)$, $\hat{a}^{j}_{\nu^{j}}(x)$, $\mu^{i},\nu^{j}=0,1...,d-1$:
@@
$\hskip-8.61108pt\bigl({l}_{0},\,{l}^{i},\,l^{ij},\,t^{ij},\,g_{0}^{i}-(s_{i}-\textstyle\frac{d}{2})\bigr)|\Phi\rangle=\bigl(\partial^{\mu}\partial_{\mu},\,-i\hat{a}^{i}_{\mu}\partial^{\mu},\,\textstyle\frac{1}{2}\hat{a}^{i}_{\mu}\hat{a}^{j\mu},\,\hat{a}^{i+}_{\mu}\hat{a}^{j\mu},\,-\hat{a}^{i+}_{\mu}\hat{a}^{\mu{}i}-s_{i}\bigr)|\Phi\rangle=\vec{0}.$ (8)
@@
The operators $o^{\prime}_{I}$ are hermitian with respect to modified (by operator $K^{\prime}$ intertwining $\{|\vec{N}\rangle_{V}\}$ and $\{|\vec{N}\rangle\}$ (30)) scalar product. Having in mind the rules of the dimensional reduction from massless theory in $\mathbb{R}^{1,d}$ space-time to massive one in $\mathbb{R}^{1,d-1}$ with fermionic oscillators $f_{i}^{+},f_{i}$, {align} ∂^M = (∂^μ, -ım) , ^a^M_i = (^a^μ_i, f_i) , ^a^M+_i = (^a^μ+_i, f_i^+) ,
M=0,1,…,d , μ=0,1,…,d-1 , η^MN = diag (1,-1,…,-1,-1) , the converted operators $O_{I}$ now for massive case are derived as well.
@@ 2.1 BRST operator and Lagrangian formulation
(for $gh_{H}(\mathcal{C}^{I})=-gh_{H}(\mathcal{P}_{I})=1$). Here, the Grassman-even ghost coordinates $q_{i}^{+}$, $q_{i}$ momenta $p_{j}$, $p_{j}^{+}$ for Grassman-odd basis elements $l_{i}$, $l_{i}^{+}$ of the superalgebra $\mathcal{A}_{C}(Y[3],\mathbb{R}^{1,d-1})$, respectively, and Grassman-odd ghost coordinates $\eta_{0}$, $\eta_{ij}^{+}$, $\eta_{ij}$, $\vartheta_{rs}^{+}$, $\vartheta_{rs}$, $\eta_{i}^{G}$ and momenta $\mathcal{P}_{0}$, $\mathcal{P}_{ij}$, $\mathcal{P}_{ij}^{+}$, $\lambda_{rs}$, $\lambda_{rs}^{+}$, $P^{G}_{i}$ for Grassman-even basis elements $l_{0}$, $L_{ij}$, $L_{ij}^{+}$, $T_{rs}$, $T_{rs}^{+}$, $G_{0}^{i}$ of $\mathcal{A}_{C}(Y[3],\mathbb{R}^{1,d-1})$, respectively, are introduced subject to non-trivial commutation relations,
@@
it follows a finite sequence of relations underlying the $\eta^{i}_{G}$-independence of all of the above homogeneous in ghost number vectors: {align} Q|χ⟩=0, (σ^i+h_i)|χ⟩=0, (ε, gh_H)(|χ⟩)=(ε_χ,0),
δ|χ⟩=Q|χ^1⟩, (σ^i+h_i)|χ^1⟩=0, (ε, gh_H)(|χ^1⟩)=(ε_χ+1,-1),
[…]
@@
As the result the operator $Q$ is nilpotent on the solutions for the middle set and the first equation in (2.1): $Q|\chi\rangle_{{s[3]}}=0$ is Lagrangian EoM and follows from gauge-invariant LF
@@
(for substitution $(K_{s[3]},Q_{s[3]})=(K,Q)|_{h^{i}=h^{i}(s[3])}$) with reducible gauge transformations {align} δ|χ^0 ⟩_s[3] =Q_s[3]|χ^1⟩_s[3] , …, δ|χ^l ⟩_s[3] =Q_s[3]|χ^l+1⟩_s[3] , l=∑_i=1^3 s_i+2 , for independent gauge parameter $|\chi^{l+1}\rangle$. Here the usual inner product for the creation and annihilation operators is assumed with measure $d^{d}x$ over Minkowski space with additional terms: "more" with auxiliary fields differed for massless and massive basic HS field $\Phi^{\mu[{s_{1}}],\nu[{s_{2}}],\rho[{s_{3}}]}(x)$.
@@ 3 Lagrangian formulation with incomplete BRST operator \\
The set of $Q_{c},\sigma^{i}_{c},\widehat{L}_{ij},\widehat{T}_{ij}$ acts in Hilbert space $\mathcal{H}^{f}\otimes H_{gh}^{o_{A}}$ and forms closed superalgebra. It provides the consistent BRST $Q_{c}$-complex resolution to get LF with holonomic operator constraints imposed on whole set of restricted field $|\chi_{c}\rangle_{s[3]}$ and reducible gauge parameter $|\chi_{c}^{n}\rangle_{s[3]}$ vectors. The latter vectors have the decomposition in oscillators like (2.1) but without conversion $b^{(+)}_{ij},d^{(+)}_{rs}$ and ghost oscillators $\eta_{ij}^{+},\mathcal{P}_{ij}^{+},\vartheta_{rs}^{+},\lambda_{rs}^{+}$ for second-class constraint system.
@@ 4 Deformation procedure for interacting vertices \\
To this end we introduce $p$, $p\geq 3$, copies of LFs (in adapting the model for Yang-Mills type interactions with gauge group $SU(N)$, for $p=N^{2}-1$) with vectors $|\chi^{(t)}_{c}\rangle_{{s}[3]_{t}}$, reducible gauge parameters $|\Lambda^{(t)n}_{c}\rangle_{{s}[3]_{t}}\equiv|\chi^{(t)n+1}_{c}\rangle_{{s}[3]_{t}}$, respective vacuum vectors $|0\rangle^{t}$ and oscillators for t=1,,pt=1,...,p (with notation $(s[3])_{t}\equiv(s[3]_{1},...,s[3]_{t})$ for arbitrary sets of indices for different copies). It permits one to determine the deformed action, gauge transformations up to $r$-tic vertices, r=3,4,,er=3,4,...,e in powers of deformation constant $g$, beginning from sum of $p$ LFs copies for free HS fields, then for cubic, quartic and so on vertices:
@@
[…]
@@
The resolution of (44) for $x$-local cubic vertices with representation